# Q-learning with Posterior Sampling

### v2

@@ 3 Algorithm design
We present our Q-learning with Posterior Sampling (PSQL) algorithm. We first present a Bayesian posterior-based derivation of the Q-learning update rule, which forms the basis for our algorithm design.
@@ 3.1 Posterior derivation
A caveat is that the above assumes $z$ to be an unbiased sample from the target distribution, whereas in Q-learning, $z$ is biased due to bootstrapping. In a recent work, Jin et al. (2018) observed that in order to account for this bias and obtain theoretical guarantees for Q-learning, the learning rate needs to be adjusted to $\alpha_{n}=\frac{H+1}{H+n}$. In fact, Bayesian inference can also provide a meaningful interpretation of this modified learning rate proposed in Jin et al. (2018). Consider the following “regularized" Bayesian inference problem Khan and Rue (2023) which adds an entropy term to the ELBO objective in (6):
@@
where ${\cal H}(q)$ denotes the entropy of the posterior. Then, we show (Lemma 7) that with weight $\lambda_{n}\mathrel{\mathop{\ordinarycolon}}=\frac{H}{n}$, when the prior $p(\theta)$ is Gaussian ${\cal N}(\widehat{\mu},\frac{\sigma^{2}}{n-1})$, and the likelihood of $z$ given $\theta$ is Gaussian ${\cal N}(\theta,\frac{\sigma^{2}}{H+1})$, then the posterior $q(\cdot)$ is given by the Gaussian distribution in (7), but with the learning rate $\alpha_{n}=\frac{H+1}{H+n}$ suggested in Jin et al. (2018). A derivation is provided in Appendix B.
Intuitively, since we suspect bias in the target $z$, (8) introduces additional uncertainty in the posterior via entropy. The weight $\lambda_{n}$ of this entropy term decreases as the number of samples increases and the bootstrapped target is expected to have lower bias. This derivation may be of independent interest as it provides a more intuitive explanation of the modified learning rate schedule proposed in Jin et al. (2018), where it was motivated mainly by the mechanics of regret analysis.
@@
The above posterior derivation forms the basis of our algorithm design presented next.
@@ 3.2 Algorithm details
$\textstyle\sigma(n)=\frac{\sigma^{2}}{n+1}\mathrel{\mathop{\ordinarycolon}}=64\frac{H^{3}}{n+1}\log(1/\delta).$ (10)
@@
and plays the arg max action
$a_{h}\mathrel{\mathop{\ordinarycolon}}=\arg\max_{a}\tilde{Q}_{h}(s_{h},a).$
The algorithm then observes a reward $r_{h}$ and the next state $s_{h+1}$, computes a target $z$, and updates the posterior mean estimate using the Q-learning update rule. A natural setting of the target would be $r_{h}+\max_{a^{\prime}}\tilde{Q}_{h+1}(s_{h+1},a^{\prime})$, which we refer to as the “vanilla version". However, due to unresolvable difficulties in regret analysis discussed later in Section 4, the PSQL algorithm computes the target in a slightly optimistic manner as we describe next. Our experiments (Appendix A) show that although this modification does impact performance, PSQL still remains significantly superior to its UCB counterparts.
1:Input: Parameters: $\delta\in(0,1)$. Set $J\mathrel{\mathop{\ordinarycolon}}=J(\delta)$ as in (11).
2:Initialize: $\widehat{Q}_{H+1}(s,a)=\widehat{V}_{H+1}(s)=0$, $\widehat{Q}_{h}(s,a)=\widehat{V}_{h}(s)=H,N_{h}(s,a)=0\,\forall s,a,h$.
3:for episodes $k=1,2,\ldots$ do
4:  Observe $s_{1}$.
5:  for step $h=1,2,\ldots,H$ do
6:   Sample $\tilde{Q}_{h}(s_{h},a)\sim\mathcal{N}(\widehat{Q}_{h}(s_{h},a),\sigma(N_{h}(s_{h},a))^{2})$, for all $a\in\mathcal{A}$.
7:   Play $a_{h}\mathrel{\mathop{\ordinarycolon}}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\tilde{Q}_{h}(s_{h},a)$.
8:   Observe $r_{h}$ and $s_{h+1}$.
9:   $z\leftarrow\texttt{ConstructTarget}(r_{h},s_{h+1},\widehat{Q}_{h+1},N_{h+1})$.
10:   $n\mathrel{\mathop{\ordinarycolon}}=N_{h}(s_{h},a_{h})\leftarrow N_{h}(s_{h},a_{h})+1$, $\alpha_{n}\mathrel{\mathop{\ordinarycolon}}=\frac{H+1}{H+n}$.
11:   $\widehat{Q}_{h}(s_{h},a_{h})\leftarrow(1-\alpha_{n})\widehat{Q}_{h}(s_{h},a_{h})+\alpha_{n}z$.
@@
1: Set $\widehat{a}=\operatorname*{arg\,max}_{a}\widehat{Q}(s^{\prime},a)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma(N(s^{\prime},a))}$.
2: /* Take maximum of the $J$ samples from the posterior of target $V_{h+1}$ */
3:Sample $\tilde{V}^{j}\sim\mathcal{N}\mathinner{\left(\widehat{Q}(s^{\prime},\widehat{a}),\sigma(N(s^{\prime},\widehat{a}))^{2}\right)}$, for $j\in[J(\delta)]$, $a\in\mathcal{A}$ .
@@
5:Return $z\mathrel{\mathop{\ordinarycolon}}=r+\overline{V}(s^{\prime})$.
Algorithm 2 ConstructTarget($r,s^{\prime},\widehat{Q},N)$
[…]
$\widehat{a}\mathrel{\mathop{\ordinarycolon}}=\arg\max_{a}\widehat{Q}_{h+1}(s^{\prime},a)+\sigma(N_{h+1}(s^{\prime},a))^{2},\text{ and}$
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\overline{V}_{h+1}(s^{\prime})=\max_{j\in J}\max_{a\in\mathcal{A}}\tilde{q}^{j}(a)$}
@@
Then,
$\overline{V}_{h+1}(s^{\prime})\mathrel{\mathop{\ordinarycolon}}=\max_{j\in[J]}\max_{a\in\mathcal{A}}\tilde{q}^{j}(a),\mbox{ and }z=r+\overline{V}(s^{\prime}).$
Observe that the above procedure computes a $\overline{V}_{h+1}(s^{\prime})$ that is more optimistic than single sample max ($\max_{a^{\prime}}\tilde{Q}_{h}(s^{\prime},a^{\prime})$). However, the optimism is limited only to the target computation not to the main decision-making in Line 6, marking an important departure from UCB based optimism (e.g., Jin et al. (2018)). The importance and rationale of this target computation procedure will be further explained in regret analysis (Section 4).
Finally, the algorithm uses the computed target $z$ to update the posterior mean via the Q-learning update rule, with $\alpha_{n}=\frac{H+1}{H+n}$:
@@ 4 Regret Analysis
Q-learning with posterior sampling poses several unique challenges for theoretical analysis.
Challenges under vanilla posterior sampling.
@@
A key difficulty in analyzing Bayesian posterior sampling algorithms is the absence of high probability optimism (i.e., the property that the estimates upper bound the true parameters). To appreciate the importance of optimism, observe that the regret of an algorithm over $K$ episodes can be decomposed into two buckets:
$\textstyle\sum_{k=1}^{K}\mathinner{\left(V^{*}_{1}(s_{k,1})-V^{\pi_{k}}_{1}(s_{k,1})\right)}\leq\underbrace{\textstyle\sum_{k=1}^{K}(V^{*}_{1}(s_{k,1})-\tilde{Q}_{k,1}(s_{k,1},a_{k,1}))}_{\text{Optimism error}}+\underbrace{\textstyle\sum_{k=1}^{K}(\tilde{Q}_{k,1}(s_{k,1},a_{k,1})-V^{\pi_{k}}_{1}(s_{k,1}))}_{\text{Estimation error}}.$ (12)
In optimistic algorithms like UCBQL Jin et al. (2018), there is no optimism error since the UCB estimate is a high-confidence upper bound on the optimal value function. For posterior sampling, in the $N$-armed bandit setting, Agrawal and Goyal (2012, 2017) made a breakthrough by showing that optimism holds with constant probability in exploration phase, and is sufficient to bound the regret. However, the recursive nature of Q-learning update makes achieving even a constant probability of optimism non-trivial in Q-learning. To see this intuitively, suppose that if we have a constant probability $p$ of optimism of posteriors on value functions in state $H$. Then, due to the recursive nature of $Q$-value update, the optimism of stage H1H-1 value requires optimism of stage $H$ value; leading to $p^{2}$ probability of optimism in stage H1H-1. Continuing this way, we get an exponentially small probability of optimism for stage $1$.
Challenges under optimistic posterior sampling.
@@
To get around the issues with constant probability optimism, many posterior sampling algorithms (e.g., model-based PSRL Agrawal and Jia (2022), MNL-bandit Agrawal et al. (2017),Tiapkin et al. (2022) etc.) taking max over multiple (say $J$) samples from the posterior in order to get high probability optimism. This modification is sometimes referred to as optimistic posterior sampling. We follow a similar modification, however, only for the target construction. We believe that, due to bootstrapping nature of Q-learning (or TD-learning methods in general) taking multiple samples for both decision-making and the target construction would lead to exponential accumulation of errors, even for $J$ as small as $2$. Below, we provide a rough intuitive argument.
@@
We expect the bias of $\tilde{Q}$ (sample from the posterior distribution) to closely track $Q^{*}$. Now, suppose $J$ samples are used in decision-making (i.e., we take multiple samples from the posterior distributions at Line 6 of Algorithm 1). Consider step $h=H$, we would incur additional regret when the bias of decision-making sample, here, $\max_{j}\tilde{Q}^{j}$ exceeds $Q^{*}$ (optimism error). Using standard techniques, this error at step $H$ has an error bound of $\sigma\sqrt{J\log(1/\delta)}$ with probability $1-\delta$ where $\sigma^{2}$ is the variance of the distribution of $\tilde{Q}$. This error subsequently propagates multiplicatively via bootstrapping of Bellman equation.
@@
For the step H1H-1, the optimism error contribution will be $\sigma\sqrt{J^{2}\log(1/\delta)}$ with probability $1-\delta$. Continuing this argument, we get that at step $1$, the cumulative optimism error will be of the order $\sigma\sqrt{J^{H}\log(1/\delta)}$, i.e., exponential in $H$ for any $J\geq 2$. To summarize, the usual trick of obtaining high probability optimism by taking multiple samples from the posterior doesn’t work for Q-learning, at least not without further novel ideas.
Our techniques.
To handle the above-mentioned challenges, our algorithm uses a combination of vanilla (single-sample) and optimistic (multiple-sample) posterior sampling for action selection and target computation, respectively. See Algorithm 1, 2 for details.
A key observation utilized here is that to break the recursive multiplicative decay in constant probability of optimism, we just need to ensure high probability optimism of the next-stage value function estimate used in the target computation. So, perhaps a solution would be to use optimistic posterior sampling in target computation only. That is, in the target computation, compute the next stage value function estimate as argmax over multiple samples from the posterior on Q-values, whereas the action taken is still decided using single sample posterior sampling. Indeed, this idea does ensure constant probability optimism while also resolving the issue of accumulating multiplicative ($J^{H}$) factors in variance and error bounds. However, now we run into the problem that the played action (single sample argmax of Q-values) may not match the action used in the target computation (multiple sample argmax). This poses difficulties in controlling the target estimation error, since the action used in the target computation may not have enough plays, and therefore, its variance may not decrease fast enough.
The final piece of the puzzle in our algorithm design and analysis is to obtain regret decomposition with only constant probability optimism for the posterior samples used for taking samples and the argmax action $\widehat{a}$ of the posterior mean (with a standard deviation offset) in our target computation. (For details, see Algorithm 2.) In our regret analysis, we show that this action $\widehat{a}$ either matches the played action with constant probability, or has a smaller variance (Lemma 5). In the next subsection, we provide a proof sketch that formally puts together all these ideas to obtain our main regret bound.

### v3

@@ 3 Algorithm design
We first present a Bayesian posterior-based derivation of the Q-learning update rule, which forms the basis for our algorithm design.
@@ 3.1 Posterior derivation
A caveat is that the above assumes $z$ to be an unbiased sample from the target distribution, whereas in Q-learning, $z$ is biased due to bootstrapping. In a recent work, Jin et al. (2018) observed that in order to account for this bias and obtain theoretical guarantees for Q-learning, the learning rate needs to be adjusted to $\alpha_{n}=\frac{H+1}{H+n}$. In fact, Bayesian inference can also provide a meaningful interpretation of this modified learning rate proposed in Jin et al. (2018). Consider the following “regularized" Bayesian inference problem (Khan and Rue, 2023) which adds an entropy term to the ELBO objective in (6):
@@
where ${\cal H}(q)$ denotes the entropy of the posterior. We show in Lemma B.1 (refer Appendix B) that for the choice of $\lambda_{n}=\frac{H}{n}$, when the prior $p(\theta)$ is Gaussian ${\cal N}(\widehat{\mu},\frac{\sigma^{2}}{n-1})$, and the likelihood of $z$ given $\theta$ is Gaussian ${\cal N}(\theta,\frac{\sigma^{2}}{H+1})$, then the posterior $q(\cdot)$ is given by the Gaussian distribution in (7) with the same learning rate $\alpha_{n}=\frac{H+1}{H+n}$ as suggested in Jin et al. (2018).
@@
The entropy regularization term of (8) introduces extra uncertainty in the posterior. Intuitively, this makes sense for Q-learning as the target $z$ is bootstrapped on previous interactions and likely has additional bias. The weight $\lambda_{n}$ of this entropy term decreases as the number of samples increases and the bootstrapped target is expected to have lower bias. This derivation may be of independent interest as it provides an intuitive explanation of the modified learning rate schedule proposed ($\frac{H+1}{H+n}$ as compared to $\frac{1}{n}$) in Jin et al. (2018), where it was motivated mainly by the mechanics of regret analysis. The above posterior derivation forms the basis of our algorithm design presented next.
@@ 3.2 Algorithm details
$\textstyle\sigma(n)=\frac{\sigma^{2}}{n+1}\mathrel{\mathop{\ordinarycolon}}=64\frac{H^{3}}{n+1}\log(KH/\delta).$ (10)
@@
and plays the arg max action $a_{h}\mathrel{\mathop{\ordinarycolon}}=\arg\max_{a}\tilde{Q}_{h}(s_{h},a)$.
The algorithm then observes a reward $r_{h}$ and the next state $s_{h+1}$, computes a target $z$, and updates the posterior mean estimate using the Q-learning update rule. A natural setting of the target would be $r_{h}+\max_{a^{\prime}}\tilde{Q}_{h+1}(s_{h+1},a^{\prime})$, which we refer to as the “vanilla version" or PSQL*. However, due to unresolvable difficulties in regret analysis discussed later in Section 4, the PSQL algorithm computes the target in a slightly optimistic manner ($r_{h}+\overline{V}_{h+1}(s)$) as we describe later in this section. Our experiments (Appendix A) show that although this modification does impact performance, PSQL still remains significantly superior to its UCB counterpart.
1:Initialize: $\widehat{Q}_{H+1}(s,a)=\widehat{V}_{H+1}(s)=0$, $\widehat{Q}_{h}(s,a)=\widehat{V}_{h}(s)=H,N_{h}(s,a)=0\,\forall s,a,h$.
2:for episodes $k=1,2,\ldots$ do
3:  Observe $s_{1}$.
4:  for step $h=1,2,\ldots,H$ do
5:   Sample $\tilde{Q}_{h}(s_{h},a)\sim\mathcal{N}(\widehat{Q}_{h}(s_{h},a),\sigma(N_{h}(s_{h},a))^{2})$, for all $a\in\mathcal{A}$.
6:   Play $a_{h}\mathrel{\mathop{\ordinarycolon}}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\tilde{Q}_{h}(s_{h},a)$.
7:   Observe $r_{h}$ and $s_{h+1}$.
8:   $z\leftarrow\texttt{ConstructTarget}(h,r_{h},s_{h+1},\widehat{Q}_{h+1},N_{h+1})$.
9:   $n\mathrel{\mathop{\ordinarycolon}}=N_{h}(s_{h},a_{h})\leftarrow N_{h}(s_{h},a_{h})+1$, $\alpha_{n}\mathrel{\mathop{\ordinarycolon}}=\frac{H+1}{H+n}$.
10:   $\widehat{Q}_{h}(s_{h},a_{h})\leftarrow(1-\alpha_{n})\widehat{Q}_{h}(s_{h},a_{h})+\alpha_{n}z$.
11:  end for
@@
Return $r$, if $h=H+1$.
 Set $\widehat{a}=\operatorname*{arg\,max}_{a}\widehat{Q}(s^{\prime},a)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma(N(s^{\prime},a))}$. Set $J\mathrel{\mathop{\ordinarycolon}}=J(\delta)$ as in (11).
/* Take maximum of the $J$ samples from the posterior of target $V_{h+1}$ */
Sample $\tilde{V}^{j}\sim\mathcal{N}\mathinner{\left(\widehat{Q}(s^{\prime},\widehat{a}),\sigma(N(s^{\prime},\widehat{a}))^{2}\right)}$, for $j\in[J(\delta)]$, $a\in\mathcal{A}$ .
$\overline{V}(s^{\prime})\leftarrow\max_{j}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{V}^{j}}$.
Return $z\mathrel{\mathop{\ordinarycolon}}=r+\overline{V}(s^{\prime})$.
Algorithm 2 ConstructTarget($h,r,s^{\prime},\widehat{Q},N)$
@@
$\widehat{a}\mathrel{\mathop{\ordinarycolon}}=\arg\max_{a}\widehat{Q}_{h+1}(s^{\prime},a)+\sigma(N_{h+1}(s^{\prime},a)),\text{ and }{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\overline{V}_{h+1}(s^{\prime})=\max_{j\in J}\tilde{V}^{j}},$
@@
Observe that the above procedure computes a $\overline{V}_{h+1}(s^{\prime})$ that is more optimistic than single sample maximum (i.e., "vanilla version" $\max_{a^{\prime}}\tilde{Q}_{h}(s^{\prime},a^{\prime})$). However, the optimism is limited only to the target computation not to the main decision-making in Line 6, marking an important departure from UCB based optimism (e.g., Jin et al. (2018)). Multiple sampling from the posteriors is a common technique considered in the past works (Tiapkin et al., 2022; Agrawal and Jia, 2017; Agrawal et al., 2017) to aid analysis. Finally, the algorithm uses the computed target $z=r_{h}+\overline{V}_{h+1}(s_{h+1})$ to update the posterior mean via the Q-learning update rule, with $\alpha_{n}=\frac{H+1}{H+n}$:
@@
Let $n=N_{k,h}(s,a)$, then Algorithm 1 implies ($\alpha_{n}^{0}\coloneqq\Pi_{j=1}^{n}(1-\alpha_{j})$ and $\alpha_{n}^{i}\coloneqq\alpha_{i}\Pi_{j=i+1}^{n}(1-\alpha_{j})$),
$\widehat{Q}_{k,h}(s,a)=\alpha_{n}^{0}H+\sum_{i=1}^{n}\alpha_{n}^{i}\mathinner{\left(r_{k_{i},h}+\overline{V}_{k_{i},h+1}(s_{k_{i},h+1})\right)}.$ (12)
@@ 4 Regret Analysis
Most of the unique challenges for the theoretical analysis of Q-learning with posterior sampling are associated with the bootstrapped nature of TD-learning itself. As shown in (12), mean estimate at the given step $h$ depends on a weighted average of the past next-step $h+1$ estimates, causing the errors at $h+1$ of the past estimates propagate to the estimate at step $h$. In model-based methods(e.g., Azar et al. (2017); Osband and Van Roy (2017); Zanette et al. (2020)) such issues are non-existent as they recalculate their estimates from scratch at each time step.
@@
Optimism dies down under recursion. One difficulty in analyzing Bayesian posterior sampling algorithms is the absence of high probability optimism (the property that the estimates upper bound the true parameters). Observe that the regret of an algorithm in any episode $k$ can be decomposed as:
$\textstyle V^{*}_{1}(s_{k,1})-V^{\pi_{k}}_{1}(s_{k,1})\leq\underbrace{\textstyle(V^{*}_{1}(s_{k,1})-\tilde{Q}_{k,1}(s_{k,1},a_{k,1}))}_{\text{Optimism error}}+\underbrace{\textstyle(\tilde{Q}_{k,1}(s_{k,1},a_{k,1})-V^{\pi_{k}}_{1}(s_{k,1}))}_{\text{Estimation error}}.$ (13)
In algorithms like UCBQL (Jin et al., 2018), there is no optimism error since the UCB estimate is a high-confidence upper bound on the optimal value function. Prior posterior sampling approaches (Agrawal and Goyal, 2012, 2017; Russo, 2019; Agrawal et al., 2021) were able to bound optimism error by proving a constant probability optimism, and then boosting to high probability by a statistical argument. However, due to the recursive nature of Q-learning, their techniques do not directly apply.
To see this, suppose that if we have a constant probability $p$ of optimism of posteriors on value functions in state $H$. The optimism of stage H1H-1 value requires optimism of stage $H$ value; leading to $p^{2}$ probability of optimism in stage H1H-1. Continuing this way, we get an exponentially small probability of optimism for stage $1$.
@@
Multiple sampling from the posteriors only partially helps. To get around the issues with constant probability optimism, many posterior sampling algorithms (e.g., model-based PSRL Agrawal and Jia (2022), MNL-bandit Agrawal et al. (2017),Tiapkin et al. (2022) etc.) taking max over multiple (say $J$) samples from the posterior in order to get high probability optimism. We follow a similar modification, with differences described later in the section. We believe that, due to bootstrapping nature of Q-learning (or TD-learning methods in general), merely taking multiple samples for either decision-making and the target construction would lead to exponential (in $H$) accumulation of errors, even for $J$ as small as $2$. Below, we provide a rough argument.
@@
We expect the bias of $\tilde{Q}_{h}$ (sample from the posterior distribution) to track $Q^{*}_{h}$ with error that scales as standard deviation of $\tilde{Q}_{h}$) (lets call that error as $\epsilon)$. Now, suppose $J$ samples are used in decision-making (i.e., we take multiple samples from the posterior distributions at Line 6 of Algorithm 1). We incur regret whenever the bias of $\max_{j}\tilde{Q}^{j}$ exceeds $Q^{*}$. Using the standard techniques, this error at step $H$ has an error bound of $\epsilon\sqrt{J\log(1/\delta)}$ with probability $1-\delta$. This error subsequently propagates multiplicatively via bootstrapping of Bellman equation. For the step H1H-1, the optimism error contribution will be $\epsilon\sqrt{J^{2}\log(1/\delta)}$ with probability $1-\delta$. Continuing this argument, at step $1$, the cumulative optimism error will be of the order $\epsilon\sqrt{J^{H}\log(1/\delta)}$, i.e., exponential in $H$ for any $J\geq 2$.
@@
The usual trick of obtaining high probability optimism by taking multiple samples from the posterior doesn’t work for Q-learning, at least not without further novel ideas.
Our techniques. The design of target computation procedure is pivotal to PSQL. Our algorithm design is characterized by two key items: (1) using optimistic posterior sampling in target computation only; (2) using the argmax action $\widehat{a}$ of the posterior mean (with a standard deviation offset) in our target computation.
First is motivated by the observation that to break the recursive multiplicative decay in constant probability of optimism, we just need to ensure high probability optimism of the next-stage value function estimate used in the target computation. Second is motivated by the previous discussion that merely taking multiple samples may lead to exponential error. In our analysis, we show that the action $\widehat{a}$ is a special action, whose standard deviation is close to the the played action $a_{h}$ with constant probability (Lemma C.2). As a result, we are able to demonstrate that as in the standard Q-learning $\overline{V}_{h}(s_{h})$ (defined with $\widehat{a}$) cannot be too far from $\tilde{Q}_{h}(s_{h},a_{h})$. Intuitively they are tracking the similar quantities. In summary, our algorithm uses a combination of vanilla (single-sample) and optimistic (multiple-sample) posterior sampling for action selection and target computation, respectively.